OpenStudy (deathslasher99):

scientific notation 123.4 x 102

OpenStudy (fibonaccichick666):

So really you have \[123.4~\times~10^2\]

OpenStudy (fibonaccichick666):

for scientific notation, it is always of the form \[a.bcdef... \times 10^2\]

OpenStudy (fibonaccichick666):

but we want it to look like \[1.234 \times 10^{?}\]

OpenStudy (fibonaccichick666):

So how id I move the decimal point

OpenStudy (deathslasher99):

2 times to the left

OpenStudy (fibonaccichick666):

\[1.234×10^{2+?}\]

OpenStudy (fibonaccichick666):

So since left is negative the ?=-2

OpenStudy (fibonaccichick666):

\[1.234\times 10^{2+-2}\]

OpenStudy (fibonaccichick666):

good so we have \[1.234\times 10^0\]

OpenStudy (fibonaccichick666):

so since everything to the zero power =1, we get \[1.234 \times 1\]

OpenStudy (fibonaccichick666):

what's that simplify to?

OpenStudy (deathslasher99):

1.234

OpenStudy (agent0smith):

@FibonacciChick666 the exponent is a positive 2, not negative

OpenStudy (fibonaccichick666):

whoops, I was trying to have an algorithm for the left right decimal movement

OpenStudy (fibonaccichick666):

whoopsie, the ?= 2

OpenStudy (fibonaccichick666):

since we are condensing so we have \[1.234 \times 10^{2+2}\]

OpenStudy (fibonaccichick666):

good so since it's 4, we get \[1.234 \times 10^4\]

OpenStudy (robtobey2):

\[123.4 * 102=12586.8=1.25868 * 10^4 \]